58.12 82.66% 44.09 18.29% PARTIV-Limiting Reactant Theoretical Yield, and Percent Yield usPoints Consider the reaction: Cu20(s) C(s) 2 Cu(s) CO(g) When 11.5 g of Care allowed to react with 114.5 g of cuzo. 87.4 g of cu are obtained.Find the limiting reactant, theoretical yield, and percent yield. Part 4
Answer #1

according to balanced reaction

12 g C reacts with 143.1 g Cu2O

11.5 g C reacts with 11.5 x 143.1 / 12 = 137.14 g Cu2O

but we have 114.5 g Cu2O

so Cu2O is limiting reagent.

143.1 gCu2O forms 2 x 63.55 g

114.5 g Cu2O forms 114.5 x 2 x 63.55 / 143.1 = 101.7 g Cu

therotical yield = 101.7 g

actual yield = 87.4 g

% yield = (87.4 / 101.7) x 100

% yield = 85.94